OpenStudy (anonymous):

Tiffany sells two kinds of homemade tomato sauce. A quart of her Tuscan sauce requires 6 tomatoes and 1 cup of oil. A quart of her marinara sauce requires 5 tomatoes and 1 1/4 cups of oil. She makes $4 profit on each quart of her Tuscan sauce and $5 profit on each quart of her marinara sauce. She has 45 tomatoes and 10 cups of oil on hand. Tiffany wants to maximize their profits from selling the sauce. Let x represent the number of quarts of Tuscan sauce and y represent the number of quarts of marinara sauce Tiffany makes. What are the constraints for the problem?

OpenStudy (anonymous):

A. 6x+y>=4 5x+5/4y>=5 x<=10 y<=10 B.6x+5y<=45 x+5/4y<=10 x>=0 y>=0 C.6x+y<=45 5x+5/4y<=10 x>=0 y>=0 D.6x+5y<=45 x+5/4y<=10 x>=4 y>=5